OpenStudy (anonymous):

Given the sequence: {2, 12, 72, 432, 2592…}, a12 = ???

OpenStudy (cwrw238):

can you see the pattern here? how do we get from 2 to 12, and then 12 to 72 by the same operation?

OpenStudy (cwrw238):

Hint - you multiply by some number

OpenStudy (cwrw238):

2 * something = 12 12 * same number = 72

OpenStudy (cwrw238):

yes and 72 * 6 = 432 and so on

OpenStudy (cwrw238):

you can write the sequence as a, ar, ar^2, ar^3 where a = 2 and r = common ratio = 6 so 12th term a12 = ar^11 = 2* 6^11

OpenStudy (cwrw238):

yes u r right